40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Mon Jul 22 16:30:31 2013, Koji, Update, General, Vent preparation
|
Mon Jul 22 16:56:14 2013, Jenne, Update, General, Vent preparation - In progress
|
Mon Jul 22 17:03:25 2013, Jenne, Update, General, MC spot positions
|
Tue Jul 23 04:26:54 2013, Jenne, Update, General, MC spot positions
|
Mon Jul 22 18:49:16 2013, Jenne, Update, General, Vent preparation - In progress
|
Mon Jul 22 21:23:04 2013, manasa, Update, General, Vent preparation - In progress
|
Tue Jul 23 03:10:57 2013, Koji, Update, General, Vent preparation - In progress 
|
Tue Jul 23 03:43:32 2013, manasa, Update, General, Vent preparation - In progress
|
Tue Jul 23 04:22:03 2013, Koji, Update, General, Vent preparation - In progress
|
Tue Jul 23 04:31:36 2013, Jenne, Update, General, Vent preparation - Ready for Vent
|
Tue Jul 23 08:21:31 2013, Steve, Update, VAC, Vent has started 
|
Tue Jul 23 14:00:03 2013, Steve, Update, VAC, the vent is completed  
|
|
Message ID: 8899
Entry time: Tue Jul 23 03:43:32 2013
In reply to: 8894
Reply to this: 8901
|
Author: |
manasa |
Type: |
Update |
Category: |
General |
Subject: |
Vent preparation - In progress |
|
|
There seems to be an unexplained oscillation in X arm cavity transmission for IR when the cavity is locked using the POX error signal.
Their origin is not related to the oplevs because the oscillation does not exist when LSC is OFF and the arms are controlled only by the oplevs and OSEMs. |